OpenStudy (anonymous):

Calculus: Find the derivative. y=cos2x/x^3

OpenStudy (anonymous):

\[y \frac{ \cos2x }{ x ^{3} }\]

OpenStudy (anonymous):

\[y=\cos2x * x ^{-3}\]

OpenStudy (anonymous):

Nw Differemtiate

OpenStudy (calculusfunctions):

\[y =\frac{ \cos 2x }{ x ^{3} }\]

OpenStudy (anonymous):

wouldn't divison rule be simpler?

OpenStudy (anonymous):

Ok...Product oR Quotient Rule...

OpenStudy (calculusfunctions):

The negative exponent is all good to use the product rule byt it is not necessary! You could use the quotient rulw! Much easier!

OpenStudy (anonymous):

ill use division rule :)

OpenStudy (calculusfunctions):

Do you know how to use the quotient rule @lilfayfay ?

OpenStudy (anonymous):

yepp but i think division rulewould be much easier :)

OpenStudy (calculusfunctions):

Same thing!!

OpenStudy (anonymous):

can we use division rule pls :)

OpenStudy (calculusfunctions):

\[\frac{ dy }{ dx }=\frac{ f \prime(x)g(x)-f(x)g \prime(x) }{ [g(x)]^{2} }\]if\[y =\frac{ f(x) }{ g(x) }\]Call it the quotient rule or call it the division rule. @lilfayfay they are the same thing!!

OpenStudy (anonymous):

oh ok lol

OpenStudy (anonymous):

so i worked out the problem and i almost got the same thing the answer is \[\frac{ -2xsin2x-3\cos2x }{ x ^{4} }\]

OpenStudy (calculusfunctions):

OK let's do this in baby steps. Step 1: Tell me the following: i). f(x) = ? ii). f '(x) = ? iii). g(x) = ? iv). g'(x) = ? v). [g(x)]² = ?

OpenStudy (calculusfunctions):

OH Is that your final answer? LOL @lilfayfay

OpenStudy (anonymous):

yeah.. :P

OpenStudy (anonymous):

sorry its late here for me kinda sleepy =w=

OpenStudy (calculusfunctions):

OH yeah your final answer is correct. I thought you were showing me the first step. Sorry.

OpenStudy (anonymous):

no, but that wasn't my final answer, that was the answer I got from the back; I got a similar answer but a little bit off so can you tell me how you did urs pls? ^^

OpenStudy (calculusfunctions):

I'm am actually a teacher LOL. I don't give out answers. You show your work and I can tell you what you did right orr start by answering my question I asked.

OpenStudy (calculusfunctions):

Why did you delete it while I was checking?

OpenStudy (calculusfunctions):

OK you really want me to write out the steps?

OpenStudy (anonymous):